Maria makes and sells exclusive scented candles. Each scented candle uses half a litre of stearin. The stearin costs SEK 100 per litre and the fragrance costs SEK 10 per candle. She then sells the candles for SEK 200 each, excluding VAT. Maria pays SEK 30,000 a month for rent and other costs.



a) Calculate the result if Maria makes and sells 4 600 candles a year.


b) Calculate the break-even point for Maria's candle production.


c) The company's current volume is 4 600 candles, calculate the margin of safety.


d) How many candles does Maria have to produce and sell in order to make a profit of SEK 300 000 per year?


1
Expert's answer
2022-02-28T11:23:01-0500

"TR=Q\u00d7P"

"=4600\u00d7200=920000"

"Profit= TR- TC"

"TC=(50+10)\u00d74600+30000=306000"

"Profit=920000-306000=614000"

"Break Even Point=Total Cost"

"=306000"